OpenStudy (jaedyn):

Solve for x: one over five (5x + 12) = 18 seventy-eight over five one hundred-two over five 6 30

OpenStudy (serenityjustice74):

so the question is 1/5(5x + 12) = 18

OpenStudy (jaedyn):

Yes

OpenStudy (serenityjustice74):

Okay. So, how do you think you should start this?

OpenStudy (serenityjustice74):

I wouldn't mind giving you the direct answer, but you aren't learning anything from this and I don't want to get both of us in trouble. So yeah.

OpenStudy (jaedyn):

I seriously dont know

OpenStudy (shadowlegendx):

\[\frac{ 1 }{ 5 } (5x + 12) = 18\]

OpenStudy (jaedyn):